OpenStudy (ziporah):

12 = −4(−6x − 3). What steps would you take to solve this for x? (distributive property?)

satellite73 (satellite73):

i wouldn't distribute, although you could

satellite73 (satellite73):

i would divide both sides by \(-4\) and start with \[-3=-6x-3\]

OpenStudy (ziporah):

If I used distributive property, would it be:? 12 = −4(−6x − 3) 12 = -4 * -6x +-4 * -3 12 = 24x + 12 12 - 12 = 24x +12 -12 0 = 24x 0/24 = 24x/24 x = 24 @satellite73

OpenStudy (ziporah):

I mean x = 0 lol

OpenStudy (kenrick143):

Yeah its still x=0 using both ways.